# Count-mean Sketch as an Optimized Framework for Frequency Estimation with Local Differential Privacy

### v1

@@ 5.2 Necessity of Randomness
[11] proposed an algorithm to convert the Hadamard Response (a special case of CMS) into a deterministic algorithm, arguing that the randomness could be avoided, thereby decreasing the communication cost. However, we will show that one can always create a counterexample causing the frequency estimation to have a constant bias. Assume that we are interested in the frequency of value $x$, and $f(x)$ is set to zero (no $x$ in the dataset). Given any hash function (denoted as $h_{j}$ for convenience), we denote $x_{j}$ as the value colliding with $x$, i.e., $c_{j}(x,x_{j})=1$. Subsequently, we set the value of all objects assigned to the $j$-th hash function as $x_{j}$, and repeat this process for all the hash functions. As a result, we have:
@@
If the assignment of hash functions to the objects, i.e., $\forall i:h_{j^{(i)}}$, is constant in CMS, one can always construct a counterexample making
@@
In Section 3.2, the result of $Var(\hat{f}(x))$ is built on $Cov(\hat{y}^{(i_{1})}[h_{j^{(i_{1})}}(x)],\hat{y}^{(i_{2})}[h_{j^{(i_{2})}}(x)])=0$ for any $i_{1}$ and $i_{2}$, which is further derived from the assumption that the hash-function assignments $\mathbf{1}(j^{(i_{1})}=j)$ and $\mathbf{1}(j^{(i_{2})}=j)$ are pairwise independent.
[…]
@@
$\lceil\max\{4\log_{2}d,2\log_{2}n\}\rceil$ random bits are enough to generate all the hash-function assignments $[j^{(i)}]_{i\in[n]}$ while ensuring their pairwise independence, which is the prerequisite of Eq. (4).
@@ 5.3 Public or Private Randomness
Previous literature (e.g., [5, 14, 12]) had a strong interest in public randomness, because it was believed that public randomness could save communication costs. However, we argue that public and private randomness make no difference for CMS, and no communication cost can be saved, for the following reasons: the server and a client (i.e., an object in CMS) have to sync on which hash function is assigned to the client, and they need to either sync (1) the identifier of a hash function, or (2) the assignment generator. There are at least $d^{2}$ hash functions in $\mathcal{H}$, so identifying one of them requires $2\log_{2}d$ bits. Meanwhile, Theorem 5.2 shows that at least $4\log_{2}d$ bits are required to represent the assignment generator. Thus, syncing the identifier will be less expensive, and $2\log_{2}d$ bits are needed to sync, regardless of which side generates the assignment. Additionally, $\log_{2}n$ bits are required if the server and a client decide to sync on the assignment generator, because the client needs to send back the object ID (i.e., $i\in[n]$) to identify itself.
@@
The only scenario where public randomness could save communication costs is when all the clients are co-located, referred to as a client group. The traffic inside a client group is assumed to be free. When utilizing the public randomness, the server will share the hash-function assignment generator with the client group, which costs $\lceil\max\{4\log_{2}d,2\log_{2}n\}\rceil$ bits. Then, the client group shares the generator, calculates their assignment of the hash functions, and replies with their responses in sequence back to the server, which will cost $n\log m\approx n\epsilon$ bits. In this case, the total communication cost between the server and the client group is $\lceil\max\{4\log_{2}d,2\log_{2}n\}\rceil+n\epsilon$. If the client group still uses private randomness, they will return $n(\log d+\epsilon)$ bits, which is larger than the cost from public randomness. However, the client-group assumption is contradictory to the motivation of LDP, which usually assumes the clients are distributed.
@@ 6 Experiment
This section empirically compares the optimized CMS with the algorithms listed below, using both synthetic and real-world datasets:
@@
Other algorithms listed in Section 5.1 are not included due to their very high communication cost, which is impractical with a very large dictionary.
@@
where $\mathbf{x}$ denotes the values of interest and $t$ denotes the experiment rounds. However, $\overline{MSE}(\hat{f})$ is biased because $\frac{1}{t}\sum_{t}(\hat{f}(x)-f(x))^{2}$ of each $x$ is equivalent to a random variable fluctuating around $E[(\hat{f}(x)-f(x))^{2}]$, and the max operator will choose the largest deviated value. Appendix I provides an upper bound of $\overline{MSE}(\hat{f})$. As long as $\overline{MSE}(\hat{f})$ is between the upper bound and the calculated $\widehat{MSE}(\hat{f})$ (both bounds are plotted), we will consider the experimental $\overline{MSE}(\hat{f})$ consistent with the theoretical value from Theorem 15.
@@
6.1 Verification of Optimality and Equivalence
@@
where $r$ is the order of a value ranked by its frequency. The dataset values are chosen deliberately, so all have the same mod for the Recursive Hadamard Response (RHR) (see Section 4.3). The dictionary size is set to 100k, and 10,000 objects are sampled from the distribution. The $\widehat{MSE}$, $l_{1}$ loss, and $l_{2}$ loss are evaluated based on the first 100 most frequent values (assuming they are known in advance). MSE-OCMS with $f^{*}=1$ (i.e., no prior knowledge on $f(x)$) and $l$-OCMS are evaluated for $\widehat{MSE}$ and $l_{1}$ / $l_{2}$ losses, respectively, while RHR, HE, and CMS+HE are evaluated for all the metrics. Results are plotted in Fig. 4.
MSE-OCMS and $l$-OCMS outperform other algorithms in terms of $\widehat{MSE}$ and $l_{1}$ / $l_{2}$ losses, respectively, and their accuracy aligns with the calculated values. HE and CMS+HE have identical performance in all the metrics, supporting our theory that these two algorithms are equivalent. Consistent with our expectation, RHR performs the worst in all the metrics because we deliberately construct a dataset, though using only public information, to ensure every object has a 50% probability of colliding with another in RHR, resulting in poorer performance than HE, as analyzed in Section 4.3.
@@ 6.2 Synthetic Dataset with Prior Knowledge
A synthetic database is constructed with 10,000 objects sampled from a normal distribution with a standard deviation of 50. The mean of the distribution is an unknown integer picked from [1000, 9000]. The dictionary size is 10,000, so the values of all the objects are rounded to integers and truncated if they are outside of [0, 10,000]. Through several simulations of this distribution, we are confident that the probability of $\max_{x}f(x)\leq 0.01$ is high. Thus, we set $f^{*}=0$ for MSE-OCMS and verify if its $\widehat{MSE}$ satisfies Eq. (16). Meanwhile, $l$-OCMS is evaluated for $l_{1}$ / $l_{2}$ losses. RHR, HE, and CMS+HE are also evaluated for $\widehat{MSE}$ $l_{1}$ loss, and $l_{2}$ loss. Results are plotted in Fig. 4.
Similar to the above, MSE-OCMS and $l$-OCMS outperform other algorithms in terms of $\widehat{MSE}$ and $l_{1}$ / $l_{2}$ losses, respectively, and their accuracy aligns with the calculated values. HE and CMS+HE have identical accuracy in all the metrics, and they perform the worst. The accuracy of RHR lies between OCMS and the other algorithms.
@@ 6.3 Real-world Dataset
Kosarak [22] is a dataset containing click-stream data from a Hungarian online news portal. The dataset is shrunk to 1% of its original size, denoted as mini Kosarak, which will be used in this section. It contains over 170,000 entries, all of which belong to a dictionary with a size of 26,000. MSE-OCMS with $f^{*}=1$ and $l$-OCMS are evaluated for $\widehat{MSE}$ and $l_{1}$ / $l_{2}$ losses, respectively, while RHR, HE, and CMS+HE are evaluated for all the metrics. Results are plotted in Fig. 4.
MSE-OCMS and $l$-OCMS outperform other algorithms in terms of $\widehat{MSE}$ and $l_{1}$ / $l_{2}$ losses, respectively, and their accuracy aligns with the calculated values. HE and CMS+HE have identical accuracy in all the metrics, and they perform the worst in $l_{1}$ / $l_{2}$ losses. RHR has the same order of accuracy as $l$-OCMS regarding $l_{1}$ / $l_{2}$ losses. However, it performs the worst in $\widehat{MSE}$. This is expected because Kosarak has one value whose frequency is above 0.4, which negatively impacts the MSE of other values sharing the same mod with it (see Section 4.3 for a detailed explanation).
@@ 7 Conclusion
This paper revisits the private Count-mean Sketch (CMS) algorithm, corrects the errors in the expectation and variance calculations in the original CMS paper, and derives the optimized parameters to reduce all kinds of loss functions. The optimized CMS is proven theoretically and empirically to be the only algorithm optimized in the worst-case MSE, $l_{1}$ loss, and $l_{2}$ loss, while its communication cost is logarithmic to the dictionary size. Moreover, this paper demonstrates that many state-of-the-art algorithms are equivalent to CMS. Finally, we discuss the necessity of randomness in CMS and show that both public and private randomness have the same communication cost.
Appendix A Derivation of the Expectation

### v2

@@ 5.2 Necessity of Randomness
An algorithm was proposed by [11] to convert the Hadamard Response (a special case of CMS) into a deterministic algorithm, arguing that the randomness could be avoided, thereby decreasing the communication cost. However, we will show that one can always create a counterexample causing the frequency estimation to have a constant bias. Assume that we are interested in the frequency of value $x$, and $f(x)$ happens to be zero (no $x$ in the dataset). Given any hash function (denoted as $h_{j}$ for convenience), we denote $x_{j}$ as the value colliding with $x$, i.e., $c_{j}(x,x_{j})=1$. Subsequently, we set the value of all objects assigned to the $j$-th hash function as $x_{j}$, and repeat this process for all the hash functions. As a result, we have:
@@
If the assignment of hash functions to the objects, $\{h_{j^{(i)}}|i\in n\}$, is constant in CMS, there exists a dataset resulting in
@@
By examining the proof involving hash-function assignments in Appendices D and E, we realize $Var(\hat{f}(x))$ is calculated based on $Cov(\hat{y}^{(i_{1})}[h_{j^{(i_{1})}}(x)],\hat{y}^{(i_{2})}[h_{j^{(i_{2})}}(x)])=0$ for any $i_{1}$ and $i_{2}$, which is further derived from the assumption that the hash-function assignments $\mathbf{1}(j^{(i_{1})}=j)$ and $\mathbf{1}(j^{(i_{2})}=j)$ are pairwise independent.
Recall that hash-function assignment uniformly samples from $\mathcal{H}$ for each object. This process maps $n$ distinct inputs (object IDs) to $n$ repeatable elements in $[k]$ (recall $k=|\mathcal{H}|$), which can be considered as hashing, denoted as $s$. We require $s$ to be uniformly sampled from a pairwise-independent hashing family $\mathcal{S}$, which guarantees the pairwise independence of $\mathbf{1}(j^{(i_{1})}=j)$ and $\mathbf{1}(j^{(i_{2})}=j)$ for all $i_{1}$, $i_{2}$, and $j$. If we follow Section 3.5 to construct a hashing family mapping from $[n]$ to $[k]$, the parameter $p^{l}$ should exceed $\max\{n,k\}$, which requires $2\lceil\log_{2}(\max\{n,k\}+1)\rceil$ bits to represent one of its mappings. We have proved:
@@
$2\lceil\log_{2}(\max\{n,k\}+1)\rceil$ random bits are sufficient to generate all the hash-function assignments $[j^{(i)}]_{i\in[n]}$ while ensuring their pairwise independence, which is the prerequisite of Eq. (4).
@@ 5.3 Public or Private Randomness
Previous literature (e.g., [5, 14, 12]) had a strong interest in public randomness, because it was believed that public randomness could save communication costs. However, we argue that public and private randomness make no difference for an unbiased CMS, and no communication cost can be saved, for the following reasons: the server and a client (i.e., an object in CMS) have to sync on which hash function is assigned to the client, and they need to either sync (1) the identifier of a hash function, or (2) the assignment generator. There are at least $(d+1)^{2}$ hash functions in a pairwise independent $\mathcal{H}_{pi}$, so identifying one of them requires $2\log_{2}d$ bits. Meanwhile, Theorem 5.2 shows that at least $2\log_{2}k$ bits are required to represent the assignment generator, where $k$ is the hashing family size, so this is equivalent to requiring at least $4\log_{2}d$ bits. Thus, syncing the identifier will be less expensive, and $2\log_{2}d$ bits are needed to sync, regardless of which side generates the assignment. Additionally, $\log_{2}n$ bits are required if the server and a client decide to sync on the assignment generator, because the client needs to send back the object ID (i.e., $i\in[n]$) to identify itself.
@@
The only scenario where public randomness could save communication costs is when all the clients are clocated, referred to as a client group. The traffic inside a client group is assumed to be free. When utilizing the public randomness, the server will share the hash-function assignment generator with the client group, which costs $\lceil\max\{4\log_{2}d,2\log_{2}n\}\rceil$ bits. Then, the client group shares the generator, calculates their assignment of the hash functions, and replies with their responses in sequence back to the server, which will cost $n\log m\approx n\epsilon$ bits. In this case, the total communication cost between the server and the client group is $\lceil\max\{4\log_{2}d,2\log_{2}n\}\rceil+n\epsilon$. If the client group still uses private randomness, they will return $n(2\log_{2}d+\epsilon)$ bits, which is larger than the cost from public randomness. However, the client-group assumption is contradictory to the motivation of LDP, which usually assumes the clients are distributed.
@@ 6 Experiment
This section empirically compares the OCMS+RR (implemented following Section 3.6) with the algorithms listed below, using both synthetic and real-world datasets:
@@
Other algorithms listed in Section 5.1 are not included due to their very high communication cost, which is impractical for a very large dictionary.
@@
where $\mathbf{x}$ denotes the values of interest and $t$ denotes the experiment rounds. However, $\overline{MSE}(\hat{f})$ is biased because $\frac{1}{t}\sum_{t}(\hat{f}(x)-f(x))^{2}$ of each $x$ is equivalent to a random variable fluctuating around $E[(\hat{f}(x)-f(x))^{2}]$, and the max operator will choose the largest deviated value. Appendix K provides an upper bound of $\overline{MSE}(\hat{f})$. As long as $\overline{MSE}(\hat{f})$ is between the upper bound and the calculated $\widehat{MSE}(\hat{f})$ (both bounds are plotted), we will consider the experimental $\overline{MSE}(\hat{f})$ consistent with the theoretical value from Theorem 13.
@@
6.1 Verification of Precision and Equivalence
@@
where $r$ is the order of a value ranked by its frequency. The dataset values are chosen deliberately, so all have the same mod for the Recursive Hadamard Response (RHR) (see Section 4.3). The dictionary size is set to 100k, and 10,000 objects are sampled from the distribution. The $\widehat{MSE}$, $l_{1}$ loss, and $l_{2}$ loss are evaluated based on the first 100 most frequent values (assuming they are known in advance). MSE-OCMS+RR with $f^{*}=1$ (i.e., no prior knowledge on $f(x)$) and $l$-OCMS+RR are evaluated for $\widehat{MSE}$ and $l_{1}$ / $l_{2}$ losses, respectively, while RHR, HE, and CMS+HE are evaluated for all the metrics. Results are plotted in Fig. 4.
MSE-OCMS+RR and $l$-OCMS+RR outperform other algorithms in terms of $\widehat{MSE}$ and $l_{1}$ / $l_{2}$ losses, respectively, and their precision aligns with the calculated values. HE and CMS+HE have identical performance in all the metrics, supporting our theory that these two algorithms are equivalent. Consistent with our expectation, RHR performs the worst in all the metrics because we deliberately construct a dataset, though using only public information, to ensure every object has a 50% probability of colliding with another in RHR, resulting in poorer performance than HE, as analyzed in Section 4.3.
@@ 6.2 Synthetic Dataset with Prior Knowledge
A synthetic database is constructed with 10,000 objects sampled from a normal distribution with a standard deviation of 50. The mean of the distribution is an unknown integer picked from [1000, 9000]. The dictionary size is 10,000, so the values of all the objects are rounded to integers and truncated if they are outside of [0, 10,000]. Through several simulations of this distribution, we are confident that the probability of $\max_{x}f(x)\leq 0.01$ is high. Thus, we set $f^{*}=0$ for MSE-OCMS+RR and its calculated $\widehat{MSE}$ is derived from Eq. (14) as:
$\widehat{MSE}(\hat{f})=\frac{1}{n}[\max_{x}f(x)+\frac{4e^{\epsilon}}{(e^{\epsilon}-1)^{2}}].$
Meanwhile, $l$-OCMS+RR is evaluated for $l_{1}$ / $l_{2}$ losses. RHR, HE, and CMS+HE are also evaluated for $\widehat{MSE}$, $l_{1}$ loss, and $l_{2}$ loss. Results are plotted in Fig. 4.
Similar to the above, MSE-OCMS+RR and $l$-OCMS+RR outperform other algorithms in terms of $\widehat{MSE}$ and $l_{1}$ / $l_{2}$ losses, respectively, and their precision aligns with the calculated values. HE and CMS+HE have identical precision in all the metrics, and they perform the worst. The precision of RHR lies between OCMS+RR and the other algorithms.
@@ 6.3 Real-world Dataset
Kosarak [23] is a dataset containing click-stream data from a Hungarian online news portal. The dataset is shrunk to 1% of its original size, denoted as mini Kosarak, which will be used in this section. It contains over 170,000 entries, all of which belong to a dictionary with a size of 26,000. MSE-OCMS+RR with $f^{*}=1$ and $l$-OCMS+RR are evaluated for $\widehat{MSE}$ and $l_{1}$ / $l_{2}$ losses, respectively, while RHR, HE, and CMS+HE are evaluated for all the metrics. Results are plotted in Fig. 4.
MSE-OCMS+RR and $l$-OCMS+RR outperform other algorithms in terms of $\widehat{MSE}$ and $l_{1}$ / $l_{2}$ losses, respectively, and their precision aligns with the calculated values. HE and CMS+HE have identical precision in all the metrics, and they perform the worst in $l_{1}$ / $l_{2}$ losses. RHR has the same order of precision as $l$-OCMS+RR regarding $l_{1}$ / $l_{2}$ losses. However, it performs the worst in $\widehat{MSE}$. This is expected because Kosarak has one value whose frequency is above 0.4, which negatively impacts the MSE of other values sharing the same mod with it (see Section 4.3 for a detailed explanation).
@@ 7 Conclusion
This paper revisits the private CMS algorithm, corrects errors in the expectation and variance calculations in the original CMS paper, and optimizes CMS with randomized response to reduce all kinds of loss functions. The optimized CMS is proven theoretically and empirically to be the leading algorithm for reducing the worst-case MSE, $l_{1}$ loss, and $l_{2}$ loss, when requiring the communication cost to be logarithmic to the dictionary size. Moreover, this paper demonstrates that many existing algorithms are equivalent to CMS. Finally, we discuss the necessity of randomness in CMS and show that both public and private randomness have the same communication cost.
Appendix A Example on Reconstruction Process
Suppose a dictionary $\mathbb{V}$ has three possible values: 1, 2, and 3; and another dictionary $\mathbb{U}$ has five possible values: $A,B,C,D$, and $E$. A mechanism $M$ perturbs a variable $V\in\mathbb{V}$ and outputs a value $U\in\mathbb{U}$. The probability that $M$ outputs $U=u$ given $V=v$ is randomly assigned as
$u$ $v$ 1 2 3
A 0.133 0.143 0.192
B 0.031 0.212 0.658
C 0.287 0.155 0.039
D 0.161 0.254 0.060
E 0.389 0.236 0.051
which is equivalent to matrix $P$. Subsequently, the inverse matrix $Q$ is derived as
$v$ $u$ A B C D E
1 0.366 -0.088 2.289 -3.629 2.267
2 0.010 -0.393 -1.928 6.625 -1.294
3 0.332 1.560 0.476 -1.996 0.240
If $M$ perturbs $V$ and outputs $U=C$, the decoding process will return the column of $Q$ corresponding to $C$, which is $[2.289,-1.928,0.476]$. If $V=1$, the probability that $M(V)$ outputs $U=C$ is 0.287. The same probability applies to the event that the reconstruction process $R(V)$, i.e., $Q[:M(V)]$, generates the corresponding column. Denote the column outputted by the reconstruction as $\tilde{V}$. One can verify that the expectation of $\tilde{V}$ equals $[1,0,0]$ when $V=1$, which is consistent with Property 2.1.
Appendix B Proof regarding Optimizing CMS with RR
We first explore the optimized $m$ for CMS + RR. Substituting Eq. (12) into Eq. (6), we have
Let us first study the worst-case MSE of CMS+RR, which is equivalent to the worst-case variance $Var(\hat{f}(x)_{RR})$. Observing that the variance is linear to $f(x)$, we have $\max_{x}MSE(\hat{f}(x))=\max\{Var(\hat{f}(x)_{RR}|f(x)=0),Var(\hat{f}(x)_{RR}|f(x)=1)\}$. When $f(x)=1$, we have
$Var(\hat{f}(x)_{RR}|f(x)=0)=Var(\hat{f}(x)_{RR}|f(x)=1),$
we have $m=e^{\epsilon/2}+1$, and the corresponding variance is
$Var(\hat{f}(x)_{RR})=\frac{e^{\epsilon/2}}{n(e^{\epsilon/2}-1)^{2}}.$
When $m\in[2,e^{\epsilon/2}+1]$, we have $Var(\hat{f}(x)_{RR}|f(x)=1)<Var(\hat{f}(x)_{RR}|f(x)=0)$. Also, $Var(\hat{f}(x)_{RR}|f(x)=0)$ decreases with $m$ in this range. Therefore,when $m\in[2,e^{\epsilon/2}+1]$, we have
That is, $m$ should never be smaller than $e^{\epsilon/2}+1$.
Define $f^{*}$ as the upper bound of $f(x)$, i.e., $\forall x:f(x)\leq f^{*}$. When $m\in[e^{\epsilon/2}+1,e^{\epsilon}+1]$, we have proven that $Var(\hat{f}(x)_{RR})$ increases with $f(x)$, so the goal becomes reducing $Var(\hat{f}(x)_{RR}|f(x)=f^{*})$. The optimized $m$ can be calculated by
$\frac{\partial Var(\hat{f}(x)_{RR}|f(x)=f^{*})}{\partial m}=0,$
which results in
$m=1+\frac{\Delta_{MSE}}{f^{*}e^{\epsilon}+1-f^{*}},$ (21)
where
The corresponding variance is
Notice that Eq. (21) is valid only when $m\in[e^{\epsilon/2}+1,e^{\epsilon}+1]$, which translates to $f^{*}\leq\frac{1}{2}$. When $f^{*}>\frac{1}{2}$, we have $m<1+e^{\epsilon/2}$, which has been proven to be worse than $m=1+e^{\epsilon/2}$. Therefore, we have
which corresponds to
Regarding Theorem 3.7, we observe that $Var(\hat{f}(x)_{RR})$ in Eq. (20) is linear to $f(x)$, so
$Var(\hat{f}(x)_{RR}|f(x)=f^{*}+\delta)-Var(\hat{f}(x)_{RR}|f(x)=f^{*})\\ =\frac{(m-1)^{2}-e^{\epsilon}}{(m-1)(e^{\epsilon}-1)n}\delta.$
Substituting $m=1+\Delta_{MSE}$ into the above equation and simplifying it, we have:
$Var(\hat{f}(x)_{RR}|f(x)=f^{*}+\delta)-Var(\hat{f}(x)_{RR}|f(x)=f^{*})\\ =\frac{\delta(1-2f^{*})e^{\epsilon}}{n\Delta_{MSE}}.$
Since $Var(\hat{f}(x)_{RR})$ increases with $f(x)$ given $m=1+\Delta_{MSE}$, we have
$\widehat{MSE}(\hat{f}(x))=Var(\hat{f}(x)_{RR}|f(x)=f^{*}+\delta)\\ =\widehat{MSE}(\hat{f}|f^{*})+\frac{\delta(1-2f^{*})e^{\epsilon}}{n\Delta_{MSE}},$
which proves the equality sign of Theorem 3.7. Given that $\frac{e^{\epsilon}}{\Delta_{MSE}}$ decreases when $f^{*}\in[0,\frac{1}{2}]$, we have $\frac{e^{\epsilon}}{\Delta_{MSE}}\leq 1$, which proves the $\leq$” sign of Theorem 3.7.
Applying RR and Substituting Eq. (12) into Eq. (22), we obtain
$l_{2}(\hat{f}_{RR})=\frac{(d+e^{\epsilon}-1)m^{2}}{n(m-1)(e^{\epsilon}-1)^{2}}+\frac{(d-1)(2m+e^{\epsilon}-1)}{n(m-1)(e^{\epsilon}-1)}.$
$m=1+\frac{\Delta_{l}}{e^{\epsilon}+d-1},$
where
In practice, we choose the closest integer to $1+\frac{\Delta_{l}}{e^{\epsilon}+d-1}$ as $m$. The corresponding optimized $l_{2}$ is the one in Theorem 3.8. Given the Cauchy-Schwarz inequality, we have $l_{1}\leq\sqrt{d\>l_{2}}$, which derives the optimized $l_{1}$ in Theorem 3.8. Corollary 3.4 is proven accordingly when applying $d\gg e^{\epsilon}$.
Appendix C Data Source of Section 5.1
We start with the precision. Since HE is equivalent to unbiased CMS+RR with $m=2$, we use Eq. (20) to compute its $\widehat{MSE}$, $l_{1}$ loss, and $l_{2}$ losses. The computed $\widehat{MSE}$ is the same as in [11]. The $l_{1}$ and $l_{2}$ losses of RHR were taken from [12], but it did not calculate the $\widehat{MSE}$. Section 4.3 shows that RHR is a skewed version of unbiased CMS, and the $\widehat{MSE}$ of RHR could be worse than that from the CMS+RR with $m=2$ (i.e., HE), so the $\widehat{MSE}$ of RHR is $\Omega\left(\left(\frac{e^{\epsilon}+1}{e^{\epsilon}-1}\right)^{2}\right)$. OLH is equivalent to the unbiased CMS with $m=1+e^{\epsilon}$, so we use Eq. (20) and follow Section 3.4 to calculate its $l_{1}$ loss, $l_{2}$ loss, and $\widehat{MSE}$. Section 4.5 shows that CMS+HE is equivalent to the unbiased CMS+RR with $m=2$ (i.e., HE), so its $l_{1}$ loss, $l_{2}$ loss, and $\widehat{MSE}$ are the same as HE. The $l_{1}$ and $l_{2}$ losses of SS are presented in [10], but their calculation of $l_{1}$ is incorrect: (1) they approximate $\sum_{x}\in[d]|\hat{f}(x)-f(x)|$ as a normal distribution, whereas the central limit theorem requires each $|\hat{f}(x)-f(x)|$ to be i.i.d, and (2) they derived the result based on the Mean Absolute Deviation of the approximated Gaussian distribution, which is equivalent to using $|\sum_{x}\hat{f}(x)-f(x)|$ to represent $\sum_{x}|\hat{f}(x)-f(x)|$. The correct way to compute $l_{1}$ is to use Cauchy–Schwarz inequality, i.e., $l_{1}\leq\sqrt{d\>l_{2}}$, which provides a tight upper bound of $l_{1}$ as $\frac{2de^{\epsilon/2}}{e^{\epsilon}-1}$. The $\widehat{MSE}$ of SS is derived using Eq. (2) in [9] with
$p=\frac{ke^{\epsilon}}{ke^{\epsilon}+d-k}$
$k=d/(e^{\epsilon}+1)$
$f=1$
The $\widehat{MSE}$ of a-RP and RP is calculated using Eq. (2) in [9] as well. Both cases where $f(x)=0$ and $f(x)=1$ are calculated, and the worst-case variance is the maximum between them.
Next, we present the data source regarding communication cost. We first compute the cost of OCMS+RR, which requires the server and a client to sync on the assigned hash function and the perturbed response $z^{(i)}$. Section 3.5 shows that $2\lceil\log_{2}\max\{d+1,5m\}\rceil$ bits are required to represent an approximately pairwise independent hash function. Also, we need another $\log_{2}m$ bits to sync the perturbed response $z^{(i)}$. Given that OCMS+RR has either $m=1+e^{\epsilon/2}$ or $m=1+e^{\epsilon}$ depending on the optimization goal, the overall communication costs of MSE-OCMS+RR and $l$-OCMS+RR are approximately $\max\{2\log_{2}d+\epsilon/2,\frac{3\epsilon}{2}+6\}$ and $\max\{2\log_{2}d+\epsilon/2,3\epsilon+6\}$, respectively.
Appendix D Derivation of the Expectation